Determine the order and inventory cost per year with the following information.

> Price of the product is $125.00 per unit.

> Market Interest is 10%

> The storage rent is $5 per unit

> The cost for the order process is $45.

> The transportation cost of an order is $100

> The quantity of the order is 24 units.

> In one year the product is ordered 25 times.

> The production cost per unit is $76.
Answer #1

Total ordering cost per year = no. of orders* ordering cost = 25*(45+100) = $3625

Total inventory cost per year = Average inventory*inventory cost per unit per year = (24/2)*5 = $60

So,

Total ordering & inventory cost per year =  3625 + 60 = $3685